shitfs --- HV problems

From: Flemming Videbaek (videbaek@sgs1.hirg.bnl.goV)
Date: Thu Sep 20 2001 - 20:04:34 EDT

  • Next message: Flemming Videbaek: "Fw: Rhc Beam Use proposal"

    The Lecroy 4032 system that maintains HV for BB counters, Tiles and the TPC gating grids i.e. all in all
    essential parts of the exp. HV went into the mode where it could not be controlled.The L1458 for TPCs, DC and RICH
    are all operational.
    Bob Scheetz and Ramiro had a controlled access around 6.00pm and decied to put the main frames on local
    after Voltages were set to proper value. This mean they cannot be controlled, nor checked by the Brahvo program. 
    The best we can hope for is that no power surge will appear and bring the PS down.
    The task of the shift crew (person) is to check that the detectors respond properly by checking the monitor plots
    carefully. Should a drop-out occur we will have to request a controlled access and reset the voltages, until we can
    attempt a diagnostic and repair phase. Unfortunately this looks like an intermittent effect, and it may well be that the systems
    are back to normal when restored to remote.
    DO NOT attempt to cycle power- this will loose settings for sure.
    contact Ramiro Debbe/F.Videbaek if system goes back in day-evening hours. At night just stand-by and note failure.
    
    
     
    
    ------------------------------------------------------
    Flemming Videbaek
    Physics Department
    Brookhaven National Laboratory
    
    tlf: 631-344-4106
    fax 631-344-1334
    e-mail: videbaek@bnl.gov
    



    This archive was generated by hypermail 2b30 : Thu Sep 20 2001 - 20:03:05 EDT